Méthodes d’analyse du cycle

1 Références centrales

Hénin [1995], Cooley [1995], Hairault [1995], Hénin [1989], Hairault [1999], King and Rebelo [1995]

2 Cycle réel

Kydland and Prescott [1982], King, Plosser, and Rebelo [1988], Cogley and Nason [1995]

3 Méthodes

Canova [1998], Kehoe and Prescott [1995]

4 Crédit, Banques, Monnaie

Carlstrom and Fuerst [1997], Christiano and Eichenbaum [1992b], Hairault and Portier [1993]

5 Macroéconomie ouverte

Obstfeld and Rogoff [1996], Obstfeld and Rogoff [1995]

6 Marché du travail

Andolfatto [1996] Merz [1995]

7 Dépenses et dette publiques

Cardia [1991] Christiano and Eichenbaum [1992a]

8 Croissance cyclique

Collard [1999]

9 Prophéties auto-réalisatrices

Farmer and Guo [1994]

10 Politiques de stabilisation

Greenwood and Huffman [1991], Hairault, Langot, and Portier [1998]

Références

Andolfatto, D. (1996) : “Business Cycle and the Labor Market Search,”

American Economic Review, 86, 112–132. Canova, F. (1998) : “Detrending and Business Cycle Facts,” Journal of

Monetary Economics, 41, 533–540. Cardia, E. (1991) : “The Dynamics of a Small Open Economy in Res-

ponse to Monetary, Fiscal and Productivity Shocks,” Journal of Monetary Economics.

Carlstrom, C., and T. Fuerst (1997) : “Agency Costs, Net Worth and Business Fluctuations : A Computable General Equilibrium Analysis,” American Economic Review, 87, 893–910.

Christiano, L., and M. Eichenbaum (1992a) : “Current Real Business Cycle Theories and Aggregate Labor-Market Fluctuations,” American Economic Review, 82(3), 430–450.

(1992b) : “Liquidity Effects and the Monetary Transmission,” American Economic Review, 82(3), 346–353.

Cogley, J., and T. Nason (1995) : “Output dynamics in Real business

Cycles Models,” American Economic Review, 85, 492–511. Collard, F. (1999) : “Spectral and Persistence Properties of Cyclical

Growth,” Journal of Economic Dynamics and control, 23, 463–488. Cooley, T. (1995) : Frontiers of Business Cycle Research. Princeton Uni-

versity Press. Farmer, R., and J. Guo (1994) : “Real Business Cycles and the Animal

Spirits Hypothesis,” Journal of Economic Theory, 63, 42–72.

Greenwood, J., and G. Huffman (1991) : “Tax Analysis in a Real Business Cycle Model : on Measuring Harberger Triangles and Okun Gaps,” Journal of Monetary Economics, 27, 167–190.

Hairault, J. (1995) : Les fluctuations conjoncturelles. Economica, Paris.

(1999) : “Vers une nouvelle synth‘ese néoclassique ?,” Revue d’Economie Politique.

Hairault, J., F. Langot, and F. Portier (1998) : “Efficiency and Stabilization : Reducing Harberger Triangles and Okun Gaps,” Miméo, M.A.D. Université de Paris I and CEPREMAP.

Hairault, J., and F. Portier (1993) : “Money, New-Keynesian Macroe-

conomics and the business Cycle,” European Economic Review. Hénin, P. (1989) : “Une macroéconomie sans monnaie pour les années 90,”

Revue d’Economie Politique, 99(4), 531–596.

(1995) : Advances in Business Cycle Research. Springer Verlag, Berlin.

Kehoe, T., and E. Prescott (1995) : “The Discipline of Applied General

Equilibrium,” Economic Theory, 6(1), 1–11. King, R., C. Plosser, and S. Rebelo (1988) : “Production, Growth and

Business Cycles I,” Journal of Monetary Economics, 21(2/3), 196–232. King, R., and S. Rebelo (1995) : “Resuscitating Real Business Cycles,” in

Handbook of Macroeconomics, ed. by R. Taylor, and M. Woodford. North-Holland, Amsterdam.

Kydland, F., and E. Prescott (1982) : “Time to Build and Aggregate

Fluctuations,” Econometrica, 50(6), 1345–1370. Merz, M. (1995) : “Search in the Labor Market and Real Business Cycle,”

Journal of Monetary Economics, 36, 269–300. Obstfeld, M., and K. Rogoff (1995) : “Exchange Rate Dynamics Re-

dux,” Journal of Political Economy, 105.

(1996) : Foundations of International Macroeconomics. The MIT Press.